OpenStudy (anonymous):

What is the x-intercept of the graph: 2x + 4y = -10

OpenStudy (anonymous):

2(x)+4y=-10 2(0)+4y=-10 4y=-10 /4 /4 y=-2.5

OpenStudy (anonymous):

Thank you!